)For a certain thermistor, its resistance at 32 o F is known to be 9 KΩ. And at 212 o F is 2KΩ. Find its resistance at 77 o F.


1
Expert's answer
2021-01-25T13:58:44-0500

Ro = 9

To  = 32 o F     is   o C + 273 = 273 o K

 RT1 2 KΩ

T1   = 212 o F   is (212 - 32) x "\frac{5}{9}"  o C + 273 = 100 + 273 = 373 o K

RT2 = ?

T2 = 77 o F      is (77- 32) x "\frac{5}{9}"  o C + 273 = 25 + 273 = 298 o K


"R_1 = R_oe^{\large \beta(\frac{1}{T_1}-\frac{1}{T_2})}"

"2 = 9e^{\large\beta(\frac{1}{373}-\frac{1}{273})}"     "ln\frac{2}{9} = \beta(\frac{1}{373}-\frac{1}{273}) \to \beta = -1.5\large\frac{373*273}{-100} =" 1531.58K

"R_2 = 2e^{1531.57(\large\frac{1}{298}-\frac{1}{373})} = 5.62K\varOmega"